OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11

OpenStudy (anonymous):

"I used substitution. Equation 1: 3x - 6y = 3 Equation 2: 7x - 5y = -11 Let's define x in terms of y in Equation 1 3x - 6y = 3 3x = 6y + 3 divide both sides by 3 x = 2y + 1 Now substitute this new definition of x into Equation 2 7x - 5y = -11 7(2y + 1) - 5y = -11 14y -5y + 7 = -11 9y = -18 divide both sides by 9 y = -2 Now substitute this value of y into our definition of x x = 1 + 2y x = 1 + 2(-2) x = 1 -4 x = -3 So in conclusion y = -2 x = -3" from yahoo lol

OpenStudy (anonymous):

haha im doing the same thing